OpenStudy (anonymous):

What is 6.4 × 10−8 mi/day in feet per hour? (1 mi = 5280 ft) A. 8.11 × 10−3 ft/h B. 1.408 × 10−5 ft/h C. 2.909 × 10−10 ft/h D. 5.05 × 10−13 ft/h

OpenStudy (anonymous):

@cc_peach

OpenStudy (anonymous):

I do these type of questions by multiplying by a series of fractions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

same

OpenStudy (anonymous):

they rn't as hard as they look

OpenStudy (anonymous):

well... w8 sometimes they r lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\frac{ 6.4 \times 10^{-8} miles }{ 1 day }\times \frac{ 5280 ft }{1 mile } \times \frac{ 1 day }{ 24 hours } = ???\frac{ ft }{ hr }\]Notice how the units cancel to end up with the units desired.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

swaggy<3

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yea im not that kools ;p

OpenStudy (anonymous):

na idk

OpenStudy (anonymous):

idc

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

did u get the answer?

OpenStudy (anonymous):

yea it is b

OpenStudy (anonymous):

Well done.

OpenStudy (anonymous):

nice job

OpenStudy (anonymous):

@cc_peach gets the medal? Boo hoo.

OpenStudy (anonymous):

lol